How much of a 5% acid solution and a 20% acid solution must be mixed in order to obtain 3 gallons of a 10% acid solution?
Please show me the steps to get the answer. Thanks

Answer by stanbon(75887)

How much of a 5% acid solution and a 20% acid solution must be mixed in order to obtain 3 gallons of a 10% acid solution?
---------------------
Equations:
Quantity Eq:: f + t = 3 gallons
Acid Eq:::: 0.05f+0.20t == 0.10*3
-----------------------------------
Multiply thru 1st equation by 5.
Multiply thru 2nd equation by 100.
------------------------------------
5f + 5t = 5*3
5f +20t = 10*3
------------------
Subtract and solve for "t":
15t = 5*3
t = 1 gallon (amt. of 20% solution needed)
---
Solve for "f":
f + t = 3
f + 1 = 3
f = 2 gallons (amt. of 5% soltuion needed)
==============================================
Cheers,
Stan H.